OpenStudy (anonymous):

PLEASE HELP... im super stuck on these 4 problems!!!! Absolute Values and Inequalities... 1. /x/+5=18 a)5 or -5 b)13 or -13 c)18 or -18 d)23 or -23 2./y+4/<1 a)-5

OpenStudy (jdoe0001):

\(\large {|x|+5=18\implies |x|=18-5\implies \begin{cases} +(x)=18-5\\ \quad \\ -(x)=18-5 \end{cases}\\ \quad \\ \quad \\ |y+4|<1\implies \begin{cases} +(y+4)<1\\ \quad \\ -(y+4)<1 \end{cases}\\ \quad \\ \quad \\ |2t|-5=7\implies |2t|=7+5\implies \begin{cases} +(2t)=7+5\\ \quad \\ -(2t)=7+5 \end{cases}\\ \quad \\ \quad \\ |a|- \cfrac{3}{4}= -\cfrac{5}{8}\implies |a|=-\cfrac{5}{8}+\cfrac{3}{4}\implies \begin{cases} +(a)=-\cfrac{5}{8}+\cfrac{3}{4}\\ \quad \\ -(a)=-\cfrac{5}{8}+\cfrac{3}{4} \end{cases}}\)

OpenStudy (anonymous):

I dont understand ...

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=wE3Us-d5KJo

OpenStudy (anonymous):

\[\left| x \right|=a,then x=-a and x=a\] \[\left| x \right|\le a ,gives -a \le x \le a\] \[\left| x \right|\ge a,gives x \le -a,x \ge a\]

OpenStudy (anonymous):

my computer blocked youtube.com

OpenStudy (anonymous):

\[\left| x+a \right|<b,gives -b <x+a <b,subtracting a from each,-b-a<x<b-a\]